OpenStudy (harlequin):

How do I set up this problem? Franklin about $12,000, part of it at the rate of 8% per year and the rest of the rate of 18% per year. If the interest received totaled $1000 how much was loaned it 8%?

OpenStudy (francesca_marie):

Try this: Let x = amount loaned at 8% and ($12,000 - x) = amount loaned at 18% Write the equation: (change the percents to decimals: 8% = 0.08 and 18% = 0.18) (0.08)x + (0.18)($12,000 - x) = $1,000

OpenStudy (francesca_marie):

This expresses the total amount of earned interest in terms of the amounts of the loan, x and ($12,000 - x}

OpenStudy (harlequin):

Franklin is supposed to be a bank loaned out

OpenStudy (francesca_marie):

Oh okay hold on

OpenStudy (harlequin):

It turned out right though so thank you

OpenStudy (francesca_marie):

Your welcome :)